OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -4) with a slope of 1/2 . y =1/2 x + 5 y =1/2 x - 3 y = 1/2x - 5 y = 1/2x + 3

OpenStudy (cwrw238):

plug in x = 2 and y = -4 into each equation and find out which one is satisfied

OpenStudy (cwrw238):

hint - the correct option will contain a minus sign ( as y = -4)

OpenStudy (cwrw238):

so you only need to check 2nd and 3rd equations

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

Could You Help Me With This One . . . Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -x + 4.

OpenStudy (anonymous):

y - 5 = (x + 1) y + 1 = (x - 5) y + 5 = -(x - 1) y - 1 = (x + 5)

OpenStudy (cwrw238):

y = -x + 4 - what is the slope of this line?

OpenStudy (cwrw238):

the slope of the following line is m: y = mx + b - the number in front of the x indicates the slope y = -x - 4 can be written y = -1x - 4 so whats the slope?

OpenStudy (anonymous):

y = -1/2x + 4

OpenStudy (cwrw238):

no the slope of y = -1x - 4 is -1 - right?

OpenStudy (anonymous):

no i was finishing the question . . . Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -1/2x + 4.